# Non-Smooth Solutions of the Navier-Stokes Equation

### v5

@@ Abstract
Non-smooth Leray-Hopf solutions of the Navier-Stokes equation are constructed. The construction occurs in a finite periodic cube $\mathbb{T}^{3}$. The entropy production maximizing solutions are selected.
Part I of this paper defines the entropy principle and using it, finds improved regularity for the Navier-Stokes solutions.
Part II concerns initial data and its achievability as a limit of the small time data.
Part III establishes analyticity properties of the Part II solution; Part IV demonstrates blowup in finite time.
I Introduction
I.1 Main Results
The main result of this paper is the construction of non-smooth entropy production maximizing solutions of the Navier-Stokes equation of the Leray-Hopf (LH) class of weak solutions.
The proof is divided into four main steps, given in Parts I, II, III, IV of this paper.
Part I formulates the fixed time entropy principle, and assuming this principle, establishes improved regularity of the Navier-Stokes solution.
Part II demonstrates the achievability of the initial data from limits of the small time data.
Part III establishes analyticity of the weak solutions of the Navier-Stokes equation.
Part IV proves blowup in finite time for initial conditions constructed with the help of Parts II and III.
@@ I.1 Main Results
I.2 Prior Results
A review of Navier-Stokes solutions can be found in BedVic22 . The criteria of BeaKatMaj84 are necessary for blowup. The numerical study Hou23 with additional numerical and theoretical references cited provide strong evidence for the existence of non-smooth solutions of the Navier-Stokes equation.
Weak solutions for the Euler equation have been constructed, with weakly defined limits and function spaces DiPMaj87 ; BreDeLSzk11 ; BreFei18 ; Neu93 ; GanWes09 ; BucVic19 ; KukVicWan20 and additional references cited in these sources. See Ber23 and references cited there for a discussion of the Navier-Stokes and Euler equations and their relation to the Onsager conjecture.
The Foias Foi74 ; FoiManRos04 ; RobRodSad16 description of Navier-Stokes turbulence as an LH weak solution provides the constructive step needed in our analysis. Other ideas, including analyticity, from this body of work are also used.
The existence of entropy production maximizing solutions of the Navier-Stokes equation is established in GliLazChe20a ; CheGliSai24b
The theory of vector and tensor spherical harmonics VarMosKhe88 is used.
The Lagrangian for space time smooth fluids, derived from Jon15 ; FukFuj10 ; FukFuj12 ; HolMarRat98 ; Tay11 ; SelWhi68 , has influenced our thinking. Scaling analysis, based on the renormalization group has also influenced our thinking. Of the many papers on this subject, we mention BogShi59 ; MonYag75 ; Wil83
@@ I.2 Prior Results
II Preliminaries
II.1 Function Spaces
@@
where $\Delta$ is the Laplace operator, FoiManRos04 , Ch. II.6, Eq. (6.16).
@@
II.2 Euler and Navier-Stokes Equations
@@
The existence of nonunique and nonphysical solutions of these equations have been known for the 1D Euler equation since the days of Riemann, for the 2D Euler equation Sch93 and in 3D even for the Navier-Stokes equation; see BucVic19 ; BucVic20 and references cited there.
@@
$\mathbb{P}_{q}$ denotes the Leray projector acting on an $L_{q}$ space, with an adjoint $\mathbb{P}_{p}$ acting on the dual $L_{p}$ space. See Eva10 .
@@
$\omega_{1}=\nabla\times u$ (5)
@@ II.3 Cutoffs
Here $\mathcal{C}^{1}_{c}$ denotes the space of continuously differentiable functions with compact support and $\mathcal{N}$ is the set of eigenfunctions of the Stokes operator, as defined in RobRodSad16 , Theorem 2.24. (Caution: $\alpha_{k}\neq a_{k}$ in (10).)
Following RobRodSad16 , Def. 3.3, a weak solution of the Navier-Stokes equation is defined as
@@
    $u$ satisfies weak conservation of energy with $\phi\in\widetilde{{D_{\sigma}}}(\mathbb{T}^{3})$ as test functions, according to RobRodSad16 , Eq. (3.1).
@@
Consequences of the definition, items 1–3, of a weak solution are given in RobRodSad16 .
By RobRodSad16 Exercise 3.4, $u$ solves the Navier-Stokes equation weakly, relative to test functions in $\mathcal{H}_{1}$. Thus $u$ solves the Navier-Stokes equation in $\mathcal{V}^{\prime}$.
By RobRodSad16 , Lemma 3.4 and 3.7,
@@
II.4 Second Order Phase Transitions of Decaying Turbulence
@@
The second order phase transition occurs in four main steps, as indicated here in a formal analysis and with the first three steps studied numerically in HsuKauGli22 , based on the analysis of DNS simulation data.
@@ II.5 Harmonic Analysis
Vector spherical harmonics describe the rotational degrees of freedom of a random vector field, such as a fluid velocity field. A possible spin variable to identify the particle transformation law is set to 1 for fluids. In this case, the vector spherical harmonics are described by three representations of $O(3)$. The full theory of tensor spherical harmonics is given in VarMosKhe88 .
II.5.1 Solid Spherical Harmonics
The customary spherical harmonics, with the basis $Y_{\ell,m}$, are augmented with a magnitude $r$. The complete expansion has the basis labeled $r,\ell,m$. On restriction to the unit sphere, $r$ takes on two values, $r=\pm 1$. The plus sign selects the solid spherical harmonics solution that is regular at the origin. The minus sign selects the solid spherical harmonics solution that is regular at infinity. Detailed accounts of spherical and solid mechanics for vector and tensor fields is given in VarMosKhe88 .
Solid mechanics harmonic analysis refers to functions that depend on $r$ as well as on the asmuthal and polar angles $\phi$ and $\theta$ which parameterize the unit sphere in $R^{3}$. MorFes53 ; Noe1918 ; WanRonBur08 ; TouSto77 . The solid mechanics solution $\Psi$, defined in terms of the parameters $(r,\theta,\phi)$ has the radial and angular separation:
@@ II.5.1 Solid Spherical Harmonics
Any function $f(r,\theta,\phi)$ has the Fourier expansion WanRonBur08 :
@@
The expansion of the total energy dissipation rate $\epsilon$ has a special form. The derivation HigKop68 shows that the index $m$ does not occur in the integral,
@@ II.5.2 Quotient Spaces
Young’s inequality FoiManRos04 , eq. (4.17), pg. 20. states that
@@
III Part I: The Entropy Principle
@@
See GliLazChe20a ; CheGliSai24b for a related discussion. Uniqueness of the entropy maximizing state will be addressed in a future publication with G.-Q. Chen and H. Said.
III.1 Dissipation Operators
We will be concerned with various forms of turbulent and viscous dissipation, as a function of $u$. The notations are given by (23).
$\begin{split}&\nu_{t,Temp}=-\langle u_{i},\partial_{t}u_{k}\rangle\\ &\nu_{t,NL}=\sum_{j}\langle u_{j}\partial_{j}u_{i}\rangle\\ &\nu_{t}=\nu_{t,NL}+\nu_{t,Temp}\\ &\nu_{Tot}=\nu_{t,Temp}+\nu_{t,NL,i},u_{k}+\nu\|\nabla u\|^{2}_{L_{s}}\ .\\ \end{split}$ (23)
Higher moments of the sums in (23) are bounded by the sum of the higher moments of the individual terms. To justify this statement, the cross terms are bounded using Young’s inequality.
Young’s inequality is used pairwise, and first serves to bound viscous vs. turbulent cross terms. The same analysis is repeated for the cross terms between the two turbulent dissipation terms.
@@ III.1 Dissipation Operators
  • Proposition III.1.1 Entropy maximizing dissipation operators $\nu_{t,Temp}$ and $\nu_{t,NL}$ are non-negative in their energy dissipation.
@@
III.2 Conservation of Energy
A weak form of the (exact) conservation of energy is contained in Lee24v6 , with an elegant and simple proof. We note the relation of this proof to earlier studies of Lad63 ; MikSoh88 ; GigSoh91 . The proof assumes a sum over repeated indices (SRI). We adapt it to the case NSRI (which is defined as no SRI). It will play a key role in the paper.
The result of Lee24 is formulated as a Lemma 1:
  • Lemma III.1 Let $u$ be a weak solution of the Navier-Stokes equation. Then
@@ III.2 Conservation of Energy
The second term vanishes by RobRodSad16 , Lemma 3.2. $\square$
@@
  • Lemma III.2 Let $u$ be an NSRI entropy production maximizing weak solution of the Navier-Stokes equation. Then the dissipation terms $\nu_{t,Temp}$ and $\nu_{t,NL}$ are bounded in $\mathcal{H}$ with norms independent of $t$.
Proof. The proof follows that of Lemma III.1 other than the assertion that the second term on the RHS of (25) is zero. The proof ends with
@@
IV Part II: Initial Data
The continuous assumption of initial data in $\mathcal{H}$ is the main issue for Part II.
To start, we eliminate a one dimensioal space of simgular solutions.
@@ IV Part II: Initial Data
  • Proposition IV.1 Let $u$ be a general solution of the Navier-Stokes equation with smooth initial data and a stirring function $f$. Then there is a solution $\tilde{u}$ and a stirring function $\tilde{f}$ both with zero global average. $\tilde{u}$ and $\tilde{f}$ are constructed from $u$ and $f$ by a global Galilean uniform drift transformation. Thus we may always assume that $u$ and $f$ have a zero spatial average.
Proof. This is a simple computation. See also FoiManRos04 , Pg. 31 and RobRodSad16 exercise 3.8. $\square$
@@
The global average is by itself a solution of the Navier-Stokes equation. It is the spherial harmonics $\ell=0$ mode of a general solution.
@@
Proof. Recall the regularity conditions (12) and item 3 of the definition of a weak solution.
The proof is identical to the proof of RobRodSad16 , Lemma 3.5. $\square$.
  • Proposition IV.2 (i) For $u$ a weak solution of the Navier-Stokes equation,
@@
    (iii) One can modify $u$ on a set of measure zero so that $u\in\mathcal{C}([0,t];\mathcal{V}^{\prime})$ and $u$ is still a solution of its transport equation.
Proof. The proof is based on RobRodSad16 , Lemmas 3.6, 3.7. The assertion (i) follows from the choice of a cutoff function $\phi$ which is constant in $t$ on the interval $[t_{1},t_{2}]$.
To prove the assertion (ii), define $g$ by the equation
$\langle g,\phi\rangle_{\mathcal{V}^{\prime}\times\mathcal{V}}=-\langle\nabla u,\nabla\phi\rangle-\langle(u\nabla)u,\phi\rangle\ .$ (29)
Then $g$ is a bounded functional on $\mathcal{V}$ for a.e. s, because
$\begin{split}&|\langle\nabla u,\nabla\phi\rangle|\leq\|\nabla u\|\|\nabla\phi\|\;\;\mathrm{and}\\ &|\langle(u\nabla u,\phi)\rangle|\leq\|(u\nabla)u\|_{L^{6/5}}\|\phi\|_{L^{6}}\\ &\leq c\|(u\nabla)u\|_{L^{6/5}}\|\nabla\phi\|\ ,\\ \end{split}$ (30)
using the Sobolev embedding $\mathcal{H}_{1}\subset L^{6}$.
It follows that $g\in L^{4/3}(0,T;\mathcal{V}^{\prime})$. Thus also $u\in L^{4/3}(0,T;\mathcal{V}^{\prime})$.
To prove assertion (iii) and also (i), we continue to follow RobRodSad16 . By choices of test functions that are constant over time intervals, we modify $u$ on a set of measure zero so that
$u\in\mathcal{C}([0,T];\mathcal{V}^{\prime})\ .$ (31)
To prove that the modified $u$ satisfies (28) for all $t_{1},t_{2}\in(0,T]$, observe that we can choose sequences $s_{n}\rightarrow t_{1}$ and $z_{n}\rightarrow t_{t}$ such that
$\begin{split}&\int_{s_{n}}^{z_{n}}\langle\nabla u,\nabla\phi\rangle+\int_{s_{n}}^{z_{n}}\langle(u\nabla)u,\phi\rangle\\ &=\langle u(s_{n}),\phi\rangle-\langle u(z_{n}),\phi\rangle\ .\\ \end{split}$ (32)
The LHS is continuous with respect to the $s_{n}$ and $z_{n}$ limits while the RHS is continuous since $u\in\mathcal{C}([0,T];\mathcal{V}^{\prime})$. In limit as $n\rightarrow\infty$, the $t_{1},t_{2}$ transport equation (28) is satisfied for all $t_{1},t_{2}>0$. The proof with the lower limit $t_{1}=0$ is similar. $\square$
@@
  • Theorem IV.1 Given a velocity $u$, there is a LH solution $u(x,t)$ of the Navier-Stokes equation with this initial condition.
@@
According to (12) or the more detailed analysis (31), the initial conditions are achieved continuously in $\mathcal{V}^{\prime}$.
The third statement is contained in the theory of FoiManRos04 . $\square$
V Part III: Analyticity
Analyticity of solutions is the central issue of Part III.

### v6

@@ Abstract
Non-smooth Leray-Hopf solutions of the Navier-Stokes equation are constructed. The construction occurs in a finite periodic cube $\mathbb{T}^{3}$. Entropy production maximizing solutions with turbulent initial data are selected. The proof of finite time blowup is based on analyticity properties of the weak solutions of the Navier-Stokes equation.
The turbulent initial data is characterized in terms of its expansion in spherical harmonics basis functions.
The mean value of a weak solution of the Navier-Stokes equation is identified as a smooth solution of the Navier-Stokes equation.
1 Introduction
1.1 Main Results
The main result of this paper is the construction of non-smooth entropy production maximizing solutions of the Navier-Stokes equation of the Leray-Hopf (LH) class of weak solutions. The initial data giving rise to the non-smooth solutions is turbulent. Turbulent initial data is identified explicitly in terms of its expansion in spherical harmonics basis elements.
Initial data which is not turbulent is characterized as mean values of general Navier-Stokes LH solutions. These solutions are entropy minimizing.
The proof is divided into three main steps.
The first step demonstrates the achievability of the initial data from limits of the small time data.
The second step establishes analyticity of the weak solutions of the Navier-Stokes equation.
The third step proves blowup in finite time for selected turbulent initial conditions.
A concluding section characterizes the mean of a weak solution of the Navier-Stokes equation as a smooth solution of the Navier-Stokes equation.
@@ 1.1 Main Results
1.2 Prior Results
A review of Navier-Stokes solutions can be found in [2]. The criteria of [1] are necessary for blowup. The numerical study [21] with additional numerical and theoretical references cited provide strong evidence for the existence of non-smooth solutions of the Navier-Stokes equation.
Weak solutions for the Euler equation have been constructed, with weakly defined limits and function spaces [10, 5, 6, 16, 7, 24] and additional references cited in these sources. See [3] and references cited there for a discussion of the Navier-Stokes and Euler equations and their relation to the Onsager conjecture.
The Foias [12, 13, 33] description of Navier-Stokes turbulence as an LH weak solution provides the constructive step needed in our analysis. Other ideas, including analyticity, from this body of work are also used.
The existence of entropy production maximizing solutions of the Navier-Stokes equation is established in [18, 9]
The theory of vector and tensor spherical harmonics [38] is used.
The Lagrangian for space time smooth fluids, derived from [23, 14, 15, 20, 36, 35], has influenced our thinking. Scaling analysis, based on the renormalization group has also influenced our thinking. Of the many papers on this subject, we mention [4, 29, 40]
@@ 1.2 Prior Results
1.3 The Fluid Millennium Prize Specification
A technical wording of this specification is given by C. Fefferman in several options. The spatially periodic option B is:
Take $\nu>0$ and $n=3$. Let $u^{0}(x)$ be any smooth, divergence-free vector field satisfying
  • smoothness, divergence free, periodicity of $u^{0}$.
  • No restrictions are placed on $f$ other than periodicity.
Then there exist smooth functions $p(x,t)$, $u_{i}(x,t)$ on $\mathbb{R}^{3}\times[0,\infty)$ that satisfy
1. the Navier-Stokes equation; 2. divergence free; 3. initial conditions given by $u^{0}(x)$; 10. periodicity; 11. smoothness.
We choose $f=0$. For initial conditions $u^{0}$ that are turbulent but otherwise unconstrained, we find a weak solution which blows up in finite time.. The prize problem is solved in the negative.
This paper and its antecedents are offered as a solution of the Millennium Fluids problem, considered as joint work of the authors contributing: J. Glimm, G.-Q. Chen, H. Said.
We show that the mean of a turbulent solution or of any weak solution is smooth. The initial data for the mean is the mean of the weak solution initial data. The mean of the weak solution data is smooth. The mean does not have general initial data. For this reason the mean is not a positive solution of the Millennium fluids problem.
2 Preliminaries
2.1 Function Spaces
@@
where $\Delta$ is the Laplace operator, [13], Ch. II.6, Eq. (6.16).
@@
2.2 Euler and Navier-Stokes Equations
@@
The existence of nonunique and nonphysical solutions of these equations have been known for the 1D Euler equation since the days of Riemann, for the 2D Euler equation [34] and in 3D even for the Navier-Stokes equation; see [7, 8] and references cited there.
@@
$\mathbb{P}_{q}$ denotes the Leray projector acting on an $L_{q}$ space, with an adjoint $\mathbb{P}_{p}$ acting on the dual $L_{p}$ space. See [11].
@@
$\omega_{i}=\nabla\times u$ (5)
@@ 2.3 Cutoffs
Here $\mathcal{C}^{1}_{c}$ denotes the space of continuously differentiable functions with compact support and $\mathcal{N}$ is the set of eigenfunctions of the Stokes operator, as defined in [33], Theorem 2.24. (Caution: $\alpha_{k}\neq a_{k}$ in (10).)
Following [33], Def. 3.3, a weak solution of the Navier-Stokes equation is defined as
@@
    $u$ satisfies weak conservation of energy with $\phi\in\widetilde{{D_{\sigma}}}(\mathbb{T}^{3})$ as test functions, according to [33], Eq. (3.1).
@@
Consequences of the definition, items 1–3, of a weak solution are given in [33].
By [33] Exercise 3.4, $u$ solves the Navier-Stokes equation weakly, relative to test functions in $\mathcal{H}_{1}$. Thus $u$ solves the Navier-Stokes equation in $\mathcal{V}^{\prime}$.
By [33], Lemma 3.4 and 3.7,
@@
2.4 Second Order Phase Transitions
@@
The second order phase transition occurs in four main steps, as indicated here in a formal analysis and with the first three steps studied numerically in [22], based on the analysis of DNS simulation data.
@@ 2.5 Harmonic Analysis
Vector spherical harmonics describe the rotational degrees of freedom of a random vector field, such as a fluid velocity field. A possible spin variable to identify the particle transformation law is set to 1 for fluids. In this case, the vector spherical harmonics are described by three representations of $O(3)$. The full theory of tensor spherical harmonics is given in [38].
2.5.1 Solid Spherical Harmonics
The customary spherical harmonics, with the basis $Y_{\ell,m}$, are augmented with a magnitude $r$. The complete expansion has the basis labeled $r,\ell,m$. On restriction to the unit sphere, $r$ takes on two values, $r=\pm 1$. The plus sign selects the solid spherical harmonics solution that is regular at the origin. The minus sign selects the solid spherical harmonics solution that is regular at infinity. Detailed accounts of spherical and solid mechanics for vector and tensor fields is given in [38].
Solid mechanics harmonic analysis refers to functions that depend on $r$ as well as on the asmuthal and polar angles $\phi$ and $\theta$ which parameterize the unit sphere in $R^{3}$. [30, 32, 39, 37]. The solid mechanics solution $\Psi$, defined in terms of the parameters $(r,\theta,\phi)$ has the radial and angular separation:
@@ 2.5.1 Solid Spherical Harmonics
Any function $f(r,\theta,\phi)$ has the Fourier expansion [39]:
@@
The expansion of the total energy dissipation rate $\epsilon$ has a special form. The derivation [19] shows that the index $m$ does not occur in the integral,
@@ 2.5.2 Quotient Spaces
Young’s inequality [13], eq. (4.17), pg. 20. states that
@@
2.6 Navier-Stokes Term by Term Regularity
If $u$ is a weak solution of the Navier-Stokes equation, then
  1. 1.
    $u\in L^{\infty}(0,T;\mathcal{H})\bigcap L^{2}(0,T;\mathcal{V})$,
  2. 2.
    $\partial_{t}u\in L^{4/3}(0,T;\mathcal{V}^{\prime})$,
  3. 3.
    $(u\cdot\nabla)u\in L^{4/3}(0,T;L^{6/5})\subset L^{4/3}(0,T;\mathcal{V}^{\prime})$,
  4. 4.
    $u\in C(0,\infty,\mathcal{V}^{\prime})$ with $C$ denoting the space of continuous functions
  5. 5.
    $\partial_{t}u-\Delta u+(u\cdot\nabla)u=0\in\mathcal{V}^{\prime}$ as an element of $L^{4/3}(0,T;\mathcal{V}^{\prime})$
by [33], Lemma 2 (with following discussion) for item 1, Lemma 3.7 for item 2, Lemma 3.4 for item 3 and Definition 3.1. Item 4 is a consequence of item 3 and item 5 is a consequence of items 1 – 3.
  • Proposition 2.6.1 The terms in Item 5 are the terms of the Navier-Stokes equation. Each term is individually spatially in $\mathcal{V}^{\prime}$ and temporally in $L^{4/3}$.
Proof. As stated above. $\square$
2.7 The Schwartz Space
The Schwartz space $\mathcal{S}$ of smooth vector fields is defined as
$\mathcal{S}=C^{\infty}(0,T;C^{\infty}(\mathbb{T}^{3}))$ (23)
with its customary topology, and its dual space is
$\mathcal{S}^{\prime}=C^{\infty^{\prime}}(0,T;C^{\infty^{\prime}}(\mathbb{T}^{3}))$ (24)
$\mathcal{S}$ and $\mathcal{S}^{\prime}$ are closed under multiplication, so that polynomials of elements of $\mathcal{S}$ are also in $\mathcal{S}$ and polynomial elements of $\mathcal{S}^{\prime}$ are also in $\mathcal{S}^{\prime}$.
2.8 Dissipation Operators
The various forms of turbulent and viscous dissipation are summarized by (25).
$\begin{split}&\nu_{t,NL}=\langle\sum_{j}u_{j}\partial_{j}u_{i}\rangle\\ &\nu_{t,Temp}=-\langle u_{i},\partial_{t}u_{k}\rangle\\ &\nu_{t}=\nu_{t,NL}+\nu_{t,Temp}\\ &\nu_{Tot}=\nu_{t,Temp}+\nu_{t,NL}+\nu\ .\\ \end{split}$ (25)
Within the solid mechanics expansion variables, each of the $r,\ell,m$ expansions are separately evaluated by the solid harmonics solution which is regular at the origin. In this way, each of the individual terms in (25) is nonnegative.
In estimating polynomials of the dissipation operators (25), cross terms are bounded by Young’s inequality, so that bounds of polynomials in the individual terms are sufficient.
3 The Entropy Principle
@@
See [18, 9] for a related discussion. Uniqueness of the entropy maximizing state will be addressed in a future publication with G.-Q. Chen and H. Said.
It is necessary to specify the physical laws assumed for the current analysis. Assuming isothermal, single fluid physics, entropy is time independent, while the energy satisfies a heat equation with decay of energy.
Other physical assumptions, such as thermal physics and compressibility allow time dependent entropy.
For the isothermal physics considered here, the entropy has a dependence on the constant temperature of the solution. A convenient normalization is to set the constant temperature to unity.
@@
  • Proposition 3.1 Entropy maximizing dissipation operators $\nu_{t,Temp}$ and $\nu_{t,NL}$ are non-negative in their energy dissipation.
@@
A weak form of the (exact) conservation of energy is contained in [26], with an elegant and simple proof. We note the relation of this proof to earlier studies of [25, 28, 17]. The proof assumes a sum over repeated indices (SRI). We adapt it to the case NSRI (which is defined as no SRI). It will play a key role in the paper.
The result of [27] is formulated as a Lemma 1:
  • Lemma 3.1 Let $u$ be a weak solution of the Navier-Stokes equation. Then
@@
The second term vanishes by [33], Lemma 3.2. $\square$
@@
  • Lemma 3.2 Let $u$ be an NSRI entropy production maximizing weak solution of the Navier-Stokes equation. Then the dissipation terms $\nu_{t,Temp}$ and $\nu_{t,NL}$ are bounded in $\mathcal{H}$ with norms independent of $t$.
Proof. The proof follows that of Lemma 3.1 other than the assertion that the second term on the RHS of (27) is zero. The proof ends with
@@
4 Initial Data
@@
  • Proposition 4.1 Let $u$ be a general solution of the Navier-Stokes equation with smooth initial data and a stirring function $f$. Then there is a solution $\tilde{u}$ and a stirring function $\tilde{f}$ both with zero global average. $\tilde{u}$ and $\tilde{f}$ are constructed from $u$ and $f$ by a global Galilean uniform drift transformation. Thus we may always assume that $u$ and $f$ have a zero spatial average.
Proof. This is a simple computation. See also [13], Pg. 31 and [33] exercise 3.8. $\square$
@@
The global average is by itself a solution of the Navier-Stokes equation. It is the spherical harmonics $\ell=0$ mode of a general solution.
@@
Proof. [33], Lemma 3.5. $\square$.
  • Proposition 4.2 (i) For $u$ a weak solution of the Navier-Stokes equation,
@@
    (iii) One can modify $u$ on a set of measure zero so that u is a continuous function of $t$ with values in $\mathcal{V}^{\prime})$ and $u$ is still a solution of its transport equation.
Proof. The proof is based on [33], Lemmas 3.6, 3.7. $\square$
@@
  • Theorem 4.1 Given a velocity $u$ with no global drift, there is a LH solution $u(x,t)$ of the Navier-Stokes equation with this initial condition.
@@
According to (12), the initial conditions are achieved continuously in $\mathcal{V}^{\prime}$.
As is conventional in the study of the Navier-Stokes equation, a global spatial average is removed by a Galilean transformation. With the subtraction of this global spatial average, the initial conditions are achieved in $\mathcal{H}$.
The third statement is contained in the theory of [13]. $\square$
5 Analyticity